# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ 2 $g(\mathcal{G}_{s})\coloneqq 0$;
4 while $openSet\neq\emptyset$ do
5    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
6    if $\mathcal{G}_{\text{current}}=\mathcal{G}_{t}$ then
7       return $g(\mathcal{G}_{t})$
8     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
    // Neighbors according to poset $\mathcal{L}$
9    foreach $\mathcal{G}_{\text{neighbor}}\in\mathrm{n}_{\mathcal{L}}(\mathcal{G}_{\text{current}}$) do
10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
11       if $g(\mathcal{G}_{\text{neighbor}})$ not defined   or   $tentative<g(\mathcal{G}_{\text{neighbor}})$ then
12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
Algorithm 1 An A* algorithm for computing the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ .
@@ 5.2 Specialization to CPDAGs
Next, we consider the concrete problem of computing $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. Here, $\mathcal{L}$ is the poset underlying all CPDAGs over a fixed vertex set. Using Algorithm 1 requires specifying a heuristic $h$ and the function $\mathrm{n}_{\mathcal{L}}$. As heuristic $h$, we use $|\mathrm{sk}(\mathcal{G})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G})|$ as a lower bound for the model-oriented distance between any $\mathcal{G}$ and $\mathcal{G}_{t}$; see Property 2 where we lower bound $m_{\text{op}}$ by zero as computing $m_{\text{op}}$ for every call of the heuristic can be costly. For the function $\mathrm{n}_{\mathcal{L}}$, we use characterization via the “Insert" and “Delete" operators introduced in Chickering (2002) (see Lemma˜2).
Computing $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ via Algorithm 1 can be expensive. We can compute the down-up and up-down distances (Property 1), to use them as upper bounds on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$. We can also compute the lower bound from Property 2. These bounds are more efficient to compute, and when they coincide, they yield the exact value of $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$. We next describe how to compute the upper bounds based on the down-up and up-down distances, and present numerical experiments comparing the model-oriented distance to the two variants of SHD. The algorithmic details for computing the lower bound are provided in Appendix I.
5.2.1 Computing an upper bound based on down-up or up-down distances
Consider first the down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$. One can adapt Algorithm 1 to search for the shortest down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. This adaptation requires only a few conceptual changes: (i) The neighbor selection step is restricted to graphs that are covered by the current graph, ensuring that each move is downward. (ii) The search terminates early when a graph $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$ is reached. At this point, the upward portion of the path is not explicitly traversed; instead, its length is computed directly as $\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$. This uses the fact that in a graded poset, the shortest upward path from $\mathcal{G}_{1}$ to $\mathcal{G}_{2}$ with $\mathcal{G}_{1}\preceq\mathcal{G}_{2}$ has length exactly $\mathrm{rank}(\mathcal{G}_{2})-\mathrm{rank}(\mathcal{G}_{1})$; see Proposition 1. (iii) The down-up distance between any graph $\mathcal{G}^{\prime}$ and $\mathcal{G}_{t}$ in this poset is given by $\mathrm{rank}(\mathcal{G}_{t})+\mathrm{rank}(\mathcal{G}^{\prime})-2\max_{\mathcal{G}\preceq\mathcal{G}^{\prime},\mathcal{G}\preceq\mathcal{G}_{t}}\mathrm{rank}(\mathcal{G})$, which motivates the following heuristic $\mathrm{rank}(\mathcal{G}_{t})+\mathrm{rank}(\mathcal{G}^{\prime})-2|\mathrm{sk}(\mathcal{G}_{t})\cap\mathrm{sk}(\mathcal{G}^{\prime})|$. In Appendix F.3, we prove that for any two CPDAGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, $\max_{\mathcal{G}\preceq\mathcal{G}_{1},\mathcal{G}\preceq\mathcal{G}_{2}}\mathrm{rank}(\mathcal{G})\leq|\mathrm{sk}(\mathcal{G}_{1})\cap\mathrm{sk}(\mathcal{G}_{2})|$, so the heuristic is a valid lower bound on the down-up distance and is therefore admissible.
Note that we check whether $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$ using the characterization in Kocka, Bouckaert and Studený (2001), which provides necessary and sufficient conditions for conditional dependence inclusion between two CPDAGs. Since this check can be computationally costly, we consider a necessary criterion for $\mathcal{G}^{\prime}$ to be the inflection point in the shortest down-up path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$. Appendix F.1 shows that the optimal inflection point is any maximizer of $\operatorname*{arg\,max}_{\mathcal{G}\preceq\mathcal{G}_{s},,\mathcal{G}\preceq\mathcal{G}_{t}}\mathrm{rank}(\mathcal{G})$, whose rank we denote by $\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})$. As previously discussed, this value is guaranteed to be less than or equal to $u=|\mathrm{sk}(\mathcal{G}_{s})\cap\mathrm{sk}(\mathcal{G}_{t})|$. Thus, before performing the check $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$, we first verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\leq\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u$. The overall adapted A* algorithm for finding the shortest down-up distance is presented in Algorithm 2.
We present the algorithm for computing the up-down distance in Appendix I, which is the ‘dual’ of Algorithm 2. In particular, (i) the neighbor selection step is restricted to be upward; (ii) the search terminates when a graph $\mathcal{G}^{\prime}$ is found with $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$; (iii) we use the admissible heuristic $2|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|-\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}^{\prime})$; and finally (iv) we verify whether $\mathrm{rank}(\mathcal{G}^{\prime})\geq|\mathrm{sk}(\mathcal{G}_{t})\cup\mathrm{sk}(\mathcal{G}^{\prime})|$ before checking if $\mathcal{G}_{t}\preceq\mathcal{G}^{\prime}$.
input: Graph class of CPDAGs $\mathfrak{G}$, pair of graphs $\mathcal{G}_{s},\mathcal{G}_{t}$, a heuristic $h\colon\mathfrak{G}\to\mathbb{R}_{\geq 0}$, downwards neighbor function $dn_{\mathcal{L}}\colon\mathfrak{G}\to 2^{\mathfrak{G}}$, upper bound $u\in\mathbb{R}_{\geq 0}$ on $\rho_{\text{meet}}(\mathcal{G}_{s},\mathcal{G}_{t})$
output : Down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
1 begin
 $openSet\coloneqq\{\mathcal{G}_{s}\}$;
 // priority queue ordered by $g+h$
2 $g(\mathcal{G}_{s})\coloneqq 0$;
3 while $openSet\neq\emptyset$ do
4    $\mathcal{G}_{\text{current}}\coloneqq$ graph in openSetopenSet minimizing $g(\mathcal{G}_{\text{current}})+h(\mathcal{G}_{\text{current}})$;
5    if $\mathrm{rank}(\mathcal{G}_{\text{current}})\leq u\,\,\text{ and }\,\,\mathcal{G}_{\text{current}}\preceq\mathcal{G}_{t}$ then
6       return $g(\mathcal{G}_{\text{current}})+\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{\text{current}})$
7     remove $\mathcal{G}_{\text{current}}$ from openSetopenSet;
8
9    foreach $\mathcal{G}_{\text{neighbor}}\in dn_{\mathcal{L}}(\mathcal{G}_{\text{current}})$ do
10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
11       if $g(\mathcal{G}_{\text{neighbor}})$ not defined   or   $tentative<g(\mathcal{G}_{\text{neighbor}})$ then
12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
Algorithm 2 An A* algorithm for computing the down-up distance $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
5.2.2 Numerical experiments for CPDAGs
To study the characteristics of the model-oriented distance $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for CPDAGs as well as the computational cost of our algorithms, we perform a small simulation study. We consider two settings. First, all pairs of CPDAGs $(\mathcal{G}_{s},\mathcal{G}_{t})$ over four nodes (185 in total; Gillispie and Perlman, 2001) and second, all pairs over five nodes ($8782$ graphs). In both settings, We first compute lower and upper bounds on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ using the procedures described earlier. When the bounds do not match, we apply Algorithm 1 to obtain the exact distance. We also compute both SHD variants (Definition 7) for comparison.
In the setting with four nodes we find the following results. For approximately $96\%$ of pairs, the bounds match exactly; for the remaining $\approx 4\%$, they differ by at most two. The average computation time is $\approx 0.00003$ seconds when bounds match, compared to $\approx 0.0001$ seconds otherwise, highlighting the efficiency of using bounds before invoking Algorithm 1. Fig.˜7(a) shows the histogram of the difference between SHD and the model-oriented distance. In $\approx 64\%$ of cases, the distances differ, and in $\approx 63\%$, the model-oriented distance is strictly smaller. Fig.˜8 presents three illustrative examples. In the first example in Figure 8(a), $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$, and thus appealing to Proposition 1, $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{rank}(\mathcal{G}_{t})-\mathrm{rank}(\mathcal{G}_{s})=2$. The SHD variants give significantly different values for the distance: $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=6$ and $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=8$. In the second example in Fig.˜8(b), we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=2$, where the shortest path between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ is an up-down path: moving a step upward from $\mathcal{G}_{s}$ to the fully connected CPDAG $\mathcal{G}$, and then moving a step downward from $\mathcal{G}$ to $\mathcal{G}_{t}$. Again, in this case, SHD variants are significantly different: $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=6$ and $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=12$. In the final example in Fig.˜8(c), $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=3$. In this case, unlike the other examples, the model-oriented distance is larger and equals $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=4$. Specifically, a shortest down-up path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ first removes the edges $v_{3}\rightarrow v_{4}$ and $v_{1}\rightarrow v_{3}$ in two downward steps to obtain a CPDAG $\mathcal{G}$ containing the edge $v_{2}-v_{4}$, and then adds the edges $v_{1}-v_{4}$ and $v_{3}-v_{4}$ in two upward steps.
In the setting with five nodes we find the following results. The upper and lower bounds on the model-oriented distance coincide in $89\%$ of cases, with an average computation time of $\approx 0.0001$ seconds. For the remaining cases, the average time is $\approx 0.002$ seconds. Figure 7(b) shows the histogram of differences between both SHD variants and the model-oriented distance. In $\approx 80\%$ of cases, the model-oriented distance differs from both SHD variants, and in $\approx 78\%$, it is strictly smaller. In about $3\%$ of cases, the shortest path follows a ‘zig-zag’ pattern rather than a down–up or up–down path. Fig.˜8(d) presents an example where the model-oriented distance and SHD differ significantly. Here, $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=10$ and $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})=12$, whereas the model-oriented distance yields $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=2$. Since SHD is model-agnostic, it does not account for the fact that all conditional dependencies encoded by $\mathcal{G}_{s}$ are also encoded by $\mathcal{G}_{t}$, and that adding two edges to $\mathcal{G}_{s}$ – one between $v_{1}$ and $v_{2}$, and another between $v_{3}$ and $v_{4}$ – yields the CPDAG $\mathcal{G}_{t}$.
Figure 7: The difference between SHD variants and our model-oriented distance, computed over all CPDAG pairs with $p$ nodes. (a) $p=4$; (b) $p=5$.
Figure 8: Examples highlighting differences between model-oriented distances and SHD; see text in Sections 5.2.2 and 5.3.1.
5.3 Specialization to MPDAGs
Next, we consider computing $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, where $\mathcal{L}_{\text{poly}}$ is the poset underlying all MPDAGs that are polytrees over a fixed vertex set. In the setting of polytrees, recall from Theorem 4 that $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ simplifies to $d_{\mathcal{L}_{\text{poly}},\text{pseudo}}(\mathcal{G}_{s},\mathcal{G}_{t})$. This simplification is important for multiple reasons: (i) appealing to Proposition 13, $\operatorname{\mathrm{SHD}}_{2}$ provides a lower bound for $d_{\mathcal{L}_{\text{poly}}}$ and thus serves as an admissible heuristic for Algorithm 1; and (ii) a graph’s neighbors in the poset $\mathcal{L}_{\text{poly}}$ are those that differ in pseudo-rank by one, and these neighbors can be computed efficiently (see Section 4.4). We supply the admissible heuristic and the neighbor function to Algorithm 1 to compute the model-oriented distance.
5.3.1 Numerical experiments for MPDAGs
We compute model-oriented distance $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$ for each pair of MPDAGs $(\mathcal{G}_{s},\mathcal{G}_{t})$ over five nodes (6679 total). The average time to compute this distance is approximately $0.004$ seconds. In roughly $98\%$ of the instances, the model-oriented distance matches the distance obtained by $\operatorname{\mathrm{SHD}}_{2}$. Among the remaining approximately $2\%$: in about $98\%$ of these cases, the model-oriented distance is two units larger than $\operatorname{\mathrm{SHD}}_{2}$; in roughly $1.7\%$ of cases, it is four units larger; and in five instances, it is six units larger. Fig.˜8(d) shows an example of an instance where the difference between the two distances is six. Here, $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})=4$, whereas the model-oriented distance yields $d_{\mathcal{L}_{\text{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=10$. The discrepancy arises because SHD incorrectly assumes that directing any of the undirected edges toward $v_{5}$ results in a valid MPDAG. However, since $\mathcal{L}_{\text{poly}}$ only contains polytrees, the shortest path in the poset $\mathcal{L}_{\text{poly}}$ from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ involves first directing each undirected edge away from $v_{5}$ (toward the other node), then removing these directed edges, and finally adding directed edges from each node to $v_{5}$.

### v2

@@ 2 $g[\mathcal{G}_{s}]\leftarrow 0$
4while $\textnormal{{openSet}}\neq\emptyset$ do
 $(\mathcal{G},\textnormal{{g}}),\textnormal{{f}}\leftarrow\textnormal{{Pop}}(\textnormal{{openSet}})$
 // Pop an item with the smallest index
5
6 if $\textnormal{{f}}\geq u^{\ast}$ then
    return $u^{\ast}$
    // Can terminate early because $\textnormal{{f}}\leq d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq u^{\ast}$
7
8
9 if $\textnormal{{g}}>g[\mathcal{G}]$ then
    Continue
    // Skip a stale entry in the queue
10
11
12 if $\mathcal{G}=\mathcal{G}_{t}$ then
    return g
    // Reached the target
13
14
15 foreach $\mathcal{G}^{\prime}\in\textnormal{{EnumNeighbors}}(\mathcal{G})$ do
    $\textnormal{{g}}^{\prime}\leftarrow\textnormal{{g}}+1$;
    // Increment current distance by one
16    $\textnormal{{u}}^{\prime}\leftarrow\textnormal{{g}}^{\prime}+u(\mathcal{G}^{\prime})$
17    if $\textnormal{{u}}^{\prime}<u^{\ast}$ then
       $u^{\ast}\leftarrow\textnormal{{u}}^{\prime}$
       // Tighten the upper bound
18
19    if $\mathcal{G}^{\prime}\notin\textnormal{{Keys}}(g)$ or $\textnormal{{g}}^{\prime}<g[\mathcal{G}^{\prime}]$ then
       $g[\mathcal{G}^{\prime}]\leftarrow\textnormal{{g}}^{\prime}$
       // Path through $\mathcal{G}^{\prime}$ is created or shortened
20       $\textnormal{{f}}^{\prime}\leftarrow\textnormal{{g}}^{\prime}+h(\mathcal{G}^{\prime})$
21       if $\textnormal{{f}}^{\prime}<u^{\ast}$ then
          $\textnormal{{Push}}(\textnormal{{openSet}},(\mathcal{G}^{\prime},\textnormal{{g}}^{\prime}),\textnormal{{f}}^{\prime})$
          // Push a potentially better solution
22
23
24
25
return 1-1
// Error: poset is not connected
Algorithm 1 Generic A* algorithm with branch and bound for computing the model-oriented distance
@@ 5.2 Specialization to CPDAGs
We specialize Algorithm˜1 to $\mathfrak{G}$, the class of CPDAGs over a vertex set $\mathrm{V}$, by specifying the subroutines required by Algorithm˜1. By Proposition˜6, an admissible heuristic can be constructed from
$d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G})\triangle\operatorname{sk}(\mathcal{G}_{t})|+2m_{\mathrm{op}}(\mathrm{S}(\mathcal{G},\mathcal{G}_{t}))\geq|\operatorname{sk}(\mathcal{G})\triangle\operatorname{sk}(\mathcal{G}_{t})|,\quad\mathcal{G}\in\mathfrak{G},$
where $|\operatorname{sk}(\mathcal{G})\triangle\operatorname{sk}(\mathcal{G}_{t})|$ and $m_{\mathrm{op}}$ can be computed using Algorithms˜2 and 3 listed in the Appendix. By the same proposition, $u(\mathcal{G})$ can be constructed from the upper bound $d_{\mathcal{L}}(\mathcal{G},\mathcal{G}_{t})\leq\min\left\{d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G},\mathcal{G}_{t}),d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G},\mathcal{G}_{t})\right\}$ for every $\mathcal{G}\in\mathfrak{G}$. Next, we describe how to compute $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G},\mathcal{G}_{t})$ and the same logic also applies to $d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G},\mathcal{G}_{t})$.
Because the poset for CPDAGs is graded, by Proposition˜3, we have
$d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G},\mathcal{G}_{t})=\operatorname{\mathrm{rank}}(\mathcal{G})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\operatorname{\mathrm{rank}}(\mathcal{G}^{\ast}),\quad\mathcal{G}\in\mathfrak{G},$
where $\operatorname{\mathrm{rank}}(\cdot)$ is given by the total number of edges (i.e., $|\operatorname{sk}(\cdot)|$). Here, $\mathcal{G}^{\ast}$ is the inflection point, i.e., a graph with maximal rank among those that are below both $\mathcal{G}$ and $\mathcal{G}_{t}$. In fact, we can use Algorithm˜1 to search for $d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G},\mathcal{G}_{t})$, only with the following adaptations: (i) $\textsc{EnumNeighbors}(\mathcal{G}^{\prime})$ is restricted to the neighbors below, i.e., those graphs covered by $\mathcal{G}^{\prime}$; (ii) the upper bound is set to
$u(\mathcal{G}^{\prime}):=\begin{cases}\operatorname{\mathrm{rank}}(\mathcal{G}_{t})+\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime}),&\quad\mathcal{G}^{\prime}\not\preceq\mathcal{G}_{t}\\ \operatorname{\mathrm{rank}}(\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime}),&\quad\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}\end{cases},$
where the first case corresponds to a path that first goes downward from $\mathcal{G}$ to $\mathcal{G}_{\hat{0}}$ and then upward to $\mathcal{G}_{t}$; (iii) the heuristic is set to
$h(\mathcal{G}^{\prime}):=\operatorname{\mathrm{rank}}(\mathcal{G}_{t})+\operatorname{\mathrm{rank}}(\mathcal{G}^{\prime})-2|\mathrm{sk}(\mathcal{G}_{t})\cap\mathrm{sk}(\mathcal{G}^{\prime})|,$
which is admissible as we show in Section˜F.2. Starting from $\mathcal{G}_{s}$, this algorithm searches downward until it finds the infection point $\mathcal{G}^{\ast}\preceq\mathcal{G}_{t}$, upon which we have $u(\mathcal{G}^{\ast})=h(\mathcal{G}^{\ast})=\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-\operatorname{\mathrm{rank}}(\mathcal{G}^{\ast})$ and the algorithm terminates immediately by line 7 of Algorithm˜1. For checking $\mathcal{G}^{\prime}\preceq\mathcal{G}_{t}$, we use the following result due to Kocka, Bouckaert and Studený (2001).
Lemma 4 (Characterization of CPDAG containment).
For CPDAGs $\mathcal{G},\mathcal{G}^{\prime}$ with model map $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}$, we have $\mathcal{G}^{\prime}\preceq\mathcal{G}$ if and only if for every pair of distinct vertices $(u,v)$,
[…]
Figure 7: Joint distribution of our distance and the structural Hamming distance computed over all pairs of CPDAGs over 4 or 5 vertices, where the identity line is drawn as dashed.
Finally, from Lemma˜2 it follows that for CPDAGs $\mathcal{G},\mathcal{G}^{\prime}$, we have $\mathcal{G}\lessdot\mathcal{G}^{\prime}$ if and only if there exist DAGs $\mathcal{D}\in[\mathcal{G}]$ and $\mathcal{D}^{\prime}\in[\mathcal{G}^{\prime}]$ such that $\mathcal{D}$ can be transformed to $\mathcal{D}^{\prime}$ through zero or more steps of covered edge reversal followed by a single edge addition. This can be turned into a subroutine that enumerates the neighbors of a CPDAG. In fact, the GES algorithm due to Chickering (2002) has already implemented this subroutine: the first phase of GES uses a set of valid Insert operations to enumerate the neighbors above (i.e., those that cover the current graph), and the second phase use a set of valid Delete operations to enumerate the neighbors below (i.e., those that are covered by the current graph). The reader is referred to Chickering (2002, §5) for a description of these operations.
Numerical results
To demonstrate our algorithm, we report results from a small numerical experiment. There are 185 CPDAGs over 4 vertices and 8782 CPDAGs over 5 vertices; see Gillispie and Perlman (2001). We use our algorithm to compute the model-oriented distance between every pair of CPDAGs over 4 or 5 vertices. Fig.˜7 compares our distance with the structural Hamming distance: both $\operatorname{\mathrm{SHD}}_{1}$ and $\operatorname{\mathrm{SHD}}_{2}$ are typically larger than the model-oriented distance. In particular, Fig.˜8 (a)-(d) highlight some pairs of graphs with large discrepancy between our distance and the $\operatorname{\mathrm{SHD}}$’s.
We implement our algorithm in C++ and the results are obtained on an ARM64 machine with 4 CPU cores and 32 GB memory. The current version scales to a moderate number of vertices. For instance, for a pair of CPDAGs over 13 vertices randomly drawn from an Erdös–Rényi model with edge probability 0.2, the algorithm takes about 0.1 second on average to compute the distance. It is worth mentioning that in this case $\mathfrak{G}$ contains around $5\times 10^{30}$ graphs (OEIS Foundation Inc., 2025), so the computation is only feasible by algorithmically exploiting the structures of the poset. From our experiments, a significant amount of computation is saved by early termination upon attaining the upper bound (line 7 of Algorithm˜1). For CPDAGs over 4 or 5 vertices, around 96% or 89% of the pairs have $d_{\mathcal{L}}=\min(d_{\mathcal{L},\uparrow\downarrow},d_{\mathcal{L},\downarrow\uparrow})$.
(a) $d_{\mathcal{L}}=2$, $\operatorname{\mathrm{SHD}}_{1}=6$, $\operatorname{\mathrm{SHD}}_{2}=8$
(b) $d_{\mathcal{L}}=2$, $\operatorname{\mathrm{SHD}}_{1}=6$, $\operatorname{\mathrm{SHD}}_{2}=12$
(c) $d_{\mathcal{L}}=4$, $\operatorname{\mathrm{SHD}}_{1}=3$, $\operatorname{\mathrm{SHD}}_{2}=3$
(d) $d_{\mathcal{L}}=2$, $\operatorname{\mathrm{SHD}}_{1}=10$, $\operatorname{\mathrm{SHD}}_{2}=12$
(e) $d_{\mathcal{L}}=10$, $\operatorname{\mathrm{SHD}}_{1}=4$, $\operatorname{\mathrm{SHD}}_{2}=4$
Figure 8: Example graph pairs that exhibit large discrepancy between the model-oriented distance and the structural Hamming distance: (a-d) are CPDAGs, (e) is polytree MPDAGs.
5.3 Specialization to polytree MPDAGs
We consider using Algorithm˜1 to compute $d_{\mathcal{L}_{\mathrm{poly}}}$ for a pair of polytree MPDAGs, where $\mathcal{L}_{\mathrm{poly}}$ is the subposet induced by polytree MPDAGs over a fixed vertex set. This is feasible because we show in Theorem˜7 that $d_{\mathcal{L}_{\mathrm{poly}}}=d_{\mathcal{L}_{\mathrm{poly}},\downarrow\uparrow}$, so we can again adapt Algorithm˜1 to compute the down-up distance. By the same theorem, the neighborhood of a graph $\mathcal{G}$ in $\mathcal{L}_{\mathrm{poly}}$ precisely consists of those graphs that differs from $\mathcal{G}$’s by unit pseudo-rank, which leads to a straightforward EnumNeighbors subroutine. Finally, $h(\mathcal{G}):=\operatorname{\mathrm{SHD}}_{2}(\mathcal{G},\mathcal{G}_{t})$ is an admissible heuristic.
Numerical results
We compute the model-oriented distance for every pair of polytree MPDAGs over 5 vertices (6679 graphs in total), which takes 0.004 second per pair on average. Fig.˜S14 in the Appendix compares our distance to the structural Hamming distances: about 98% of the pairs have $d_{\mathcal{L}_{\mathrm{poly}}}=\operatorname{\mathrm{SHD}}_{2}$. Fig.˜8(d) shows an example where $d_{\mathcal{L}_{\mathrm{poly}}}=\operatorname{\mathrm{SHD}}_{2}+6$.